AS6502(1L) AS6502(1L) NAME as6502,as65c00,asr65c00−crossassemblersformicrocomputers SYNOPSIS as6502[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as65c00[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input asr65c00[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas6502,as65c00,asr65c00commandassemblestheinputfileintoatextoutputfilerepresentingtheprogrammemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu 65,65c,r65,r65c,21 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas6502Frankenstein Assembler(fileas6502.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch.Finallytheinputfilecan selectwhichsubsettousewiththeCPUstatement. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1